# A Brain-Inspired Regularizer for Adversarial Robustness

### v1

@@ 5.6 Computational advantages
In addition to being a simple method to apply, our regularization method is computationally inexpensive. First, in regard to training time, for $k$ image pairs per regularize batch, the additional time taken per batch to train the model corresponds to $2\times k$ additional forward passes. We see in Fig. 8 that the method is successful for regularization batch size values: 4, 8, 16, 32. Choosing a smaller batch size can help in cutting the extra training time needed for successful regularization.
@@
Second, although the number of target similarities is $\binom{N}{2}$ for $N$ selected regularization images, we find that we do not need many images for regularization. In our experiments we used $N=5000$, leading to approx $12\times 10^{6}$ pairs (Appendix 7.1), however, we show in Fig. 9 that we do not need that much images; as can be seen using only $N\in\{100,1000\}$ images provides robustness increase to black box attacks. Last, our method relies on the original image datasets, and does not require the introduction of different data distortions or augmentations during training.
@@ 5.7 Results using color datasets
Figure 10: Robustness of a ResNet18 trained to classify colored CIFAR-10 regularized on colored images from different datasets : CIFAR-10 (blue), CIFAR-100 (purple) or ImageNet (red). For the decision-based Boundary Attack, we compute the median $L_{2}$ perturbation size, averaged across 1000 images, and 5 repeats. Error shades/bars represent the SEM across seven seeds per model.
@@ 6 Conclusion and discussion
We showed that our proposed method increases the robustness of CNNs to a spectrum of black box attacks (Section 2). We proposed a method to select the regularization hyperparameters ($\alpha,Th$). We also showed that the choice of an ($\alpha$, $Th$) pair value for regularization, affects the robustness level in a consistent way across different attacks. We demonstrated the effectiveness and scalability of our method, by showing its success in increasing model robustness using different combinations of classification and regularization datasets, including classification datasets CIFAR-10 and CIFAR-100 (Appendix 7.1). We evaluated the performance of regularized models on common corruptions using grayscale CIFAR-10-C. We performed a Fourier analysis on minimal adversarial perturbations obtained from a decision-based Boundary Attack on our regularized model, and found that the perturbations from the regularized model contained higher low-frequency components relative to the unregularized model. We also showed that our method is more effective against common corruptions that are categorized as high frequency corruptions based on the average frequency estimated from the Fourier spectrum of the perturbations induced by these corruptions [6]. These findings are in line to those in Li et al. [6], who examined the same perturbations for a model regularized using neural data [1]. Even though we mostly presented results using grayscale datasets to allow direct comparison with the method in Li et al. [1], we demonstrated that our method is also successful when using color datasets, where we showed results for color CIFAR-10 and color CIFAR-100. We also investigated the contribution of different parts of $S^{Th}$ (Appendix 7.7).

### v2

@@ 5.6 Computational advantages
In addition to being a simple method to apply, our regularization method is computationally inexpensive. First, in regard to training time, for $k$ image pairs per regularization batch, the additional time taken per batch to train the model corresponds to $2\times k$ additional forward passes. We see in Fig. 8 that the method is successful for regularization batch size values: 4, 8, 16, 32. Choosing a smaller batch size can help in cutting the extra training time needed for successful regularization.
@@
Second, although the number of target similarities is $\binom{N}{2}$ for $N$ selected regularization images, we find that we do not need many images for regularization. In our experiments we used $N=5000$, leading to approximately $12\times 10^{6}$ pairs (Appendix 7.1), however, we show in Fig. 9 that we do not need that much images; as can be seen using only $N\in\{100,1000\}$ images provides robustness increase to black box attacks. Last, our method relies on the original image datasets, and does not require the introduction of different data distortions or augmentations during training.
@@ 5.7 Results using color datasets
Figure 10: Robustness of a ResNet18 trained to classify color CIFAR-10 regularized on colored images from different datasets : CIFAR-10 (blue), CIFAR-100 (purple) or ImageNet (red). For the decision-based Boundary Attack, we compute the median $L_{2}$ perturbation size, averaged across 1000 images, and 5 repeats. Error shades/bars represent the SEM across seven seeds per model.
@@ 6 Conclusion and discussion
We showed that our method increases the robustness of CNNs to a spectrum of black box attacks (Section 2). We proposed a method to select the regularization hyperparameters ($\alpha,Th$), and showed that the choice of an ($\alpha$, $Th$) pair value for regularization, affects the robustness level in a consistent way across different attacks. We demonstrated the effectiveness and scalability of our method, by showing its success in increasing model robustness using different combinations of classification and regularization datasets, including classification datasets CIFAR-10 and CIFAR-100 (Appendix 7.1). We evaluated the performance of regularized models on common corruptions using grayscale CIFAR-10-C. We performed a Fourier analysis on minimal adversarial perturbations obtained from a decision-based Boundary Attack on our regularized model, and found that the perturbations from the regularized model contained higher low-frequency components relative to the unregularized model. We also showed that our method is more effective against common corruptions that are categorized as high frequency corruptions based on the average frequency estimated from the Fourier spectrum of the perturbations induced by these corruptions [6]. These findings are in line to those in Li et al. [6], who examined the same perturbations for a model regularized using neural data [1]. Even though we mostly presented results using grayscale datasets to allow direct comparison with the method in Li et al. [1], we demonstrated that our method is also successful when using color datasets, where we showed results for color CIFAR-10 and color CIFAR-100. We also investigated the contribution of different parts of $S^{Th}$ (Appendix 7.7).